A bar with a square cross section with dimension 14 mm x 14 mm has a force P = 15 kN applied to it axially. The length L of the bar is 497 mm and has a Young’s modulus of 94 GPa. Calculate the extension of the bar due to the axial load applied. Express your answer in mm.


1
Expert's answer
2021-02-17T10:41:53-0500

Find the stress:


"\sigma=\frac FA."

Express the strain in terms of stress and Young's modulus:


"\epsilon=\frac{\sigma}{E}=\frac{F}{AE}."


Find the extension:


"\Delta L=L\epsilon=\frac{FL}{AE}=0.4\text{ mm}."